OpenStudy (anonymous):

How do you find intercepts of 8x=12y-18

OpenStudy (anonymous):

X-intercept: Let y = 0, and solve for x Y-intercept: Let x = 0, and solve for y Now, get to work!

OpenStudy (anonymous):

I wouldn't need to move anythig around?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

Follow what I showed you above.

OpenStudy (anonymous):

Ah. Okay. Thanks.

OpenStudy (anonymous):

For the x-intercept, let y = 0, so you get 8x = 12(0) - 18 8x = -18 x = -18/8 or -9/4 Thats the x-intercept. See how easy that was.

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

For the y-intercept, let x = 0 So, 8(0) = 12 y - 18 0 = 12y - 18 12y = 18 y = 1.5 or 3/2 The y-intercept is 3/2 or 1 1/2. So simple.

OpenStudy (anonymous):

By the way, now that you have the x-intercept and y-intercepot, you can draw the line by plotting these 2 points and connect the straight line.

OpenStudy (anonymous):

Now, one more thing.

OpenStudy (anonymous):

I got that part haha. Thanks. :3

OpenStudy (anonymous):

For the y-intercept, we let x = 0. Why? Because the y-intercept is the point on the y-axis. And any point on the y-axis, x must be 0. Same idea for the x-intercept; every point on the x-axis, y must be zero. Mathematics makes sense!

OpenStudy (anonymous):

Good luck!